Find the first derivative of y with respect to x. Use the given relation in its

implicit form.

a. x2 + y2 = a2

b. X2 + 4y2 = 4ay


1
Expert's answer
2022-03-02T17:02:16-0500

"a. x^2 +y^2=a^2\\
2x+2yy'=0\\
y'=-\frac{2x}{2y}\\
y'=-\frac{x}{y}\\
b. x^2 +4y^2=4ay\\
2x+8yy'=4ay'\\
4ay'-8yy'=2x\\
y'(4a-8y)=2x\\
y'=\frac{2x}{4a-8y}\\
y'=\frac{x}{2a-4y}\\"